# Scalable Unit Harmonization in Medical Informatics via Bayesian-Optimized Retrieval and Transformer-Based Re-ranking

### v2

@@ 1 Background and Significance
Modern scientific research increasingly depends on the integration of heterogeneous data sources, with inconsistent units of measurement posing a persistent challenge, particularly in laboratory analyses where identical quantities may be reported using diverse conventions. Data harmonization is a critical prerequisite for large-scale clinical research, ensuring interoperability across electronic health record systems and supporting reliable, reproducible medical analyses. Laboratory unit harmonization is especially complex due to domain-specific terminology and the clinical importance of precise unit conversion. Traditional methods for unit harmonization often rely on manually curated rule-based systems and mapping tables developed by domain experts [1]. While these approaches can be effective within narrowly defined contexts, they are not easily scalable and require ongoing maintenance to accommodate emerging laboratory codes and units. This labor-intensive process introduces bottlenecks in research workflows, delays analysis, and increases resource burdens. Emerging methodologies that combine machine learning with information retrieval techniques offer promising solutions to these challenges. Techniques such as sentence embeddings, bidirectional transformers, and optimized retrieval models have demonstrated potential for automating and scaling harmonization processes with improved accuracy and efficiency [2, 3, 4]. In this paper, we present a novel automated unit harmonization system that advances current practices by integrating Bayesian-optimized BM25, sentence embeddings, and transformer-based re-ranking mechanisms. Our system directly addresses major limitations in existing frameworks, including poor scalability to billions of records, limited adaptability to diverse naming conventions, inadequate contextual understanding of laboratory codes, lack of dynamic feedback loops, and overreliance on manual validation. By mitigating these constraints, our approach facilitates seamless data integration, reduces error rates, and accelerates the pace of scientific discovery, ultimately empowering clinical researchers to derive more robust, timely, and reproducible insights from increasingly complex and voluminous healthcare datasets.
@@ 2.2.1 Overall Pipeline Structure
The system is organized as an end-to-end processing pipeline (Fig. 1) comprising three main stages: preprocessing, harmonization, and post-processing.
The preprocessing stage begins by retrieving raw data from the source system, Optum Clinformatics Data Mart (2024 Q1, Q2, and Q3). To support downstream harmonization, key fields are extracted, including LOINC codes, units of measurement, frequency of occurrence, and descriptive statistics (minimum, maximum, mean, and standard deviation). This process yields approximately 30,000 unique (test, sample, unit) triads requiring standardization, of which 17,244 were found to have candidate matches in our internal database.
Next, the extracted data undergoes cleaning and normalization to ensure consistency in structure and values. A validation step then confirms the integrity of the LOINC codes and associated metadata. Finally, the dataset is enriched by mapping valid LOINC codes to standardized labels and attributes, providing essential context for the harmonization phase.
@@
During harmonization, the system constructs tailored queries that generate ranked proposals for matching equivalents. These queries leverage the hybrid retrieval engine presented in this paper, enabling efficient searches for equivalent data elements. The retrieved results are then processed and structured for evaluation, followed by a reranking step that prioritizes candidates based on relevance scores to enhance overall accuracy.
@@ 2.2.2 Predictive Model Architecture
The Retrieval Reranker serves as the final decision layer. This bidirectional transformer model [6, 7, 8] reevaluates and reorders candidate matches based on deeper contextual understanding, addressing subtle semantic distinctions critical in complex medical terminology.
@@ 2.2.3 Baseline Retrieval Using BM25
To further enhance retrieval effectiveness, the system incorporates curated synonym lists that expand the query space by mapping equivalent or related terms. This extension is particularly valuable in laboratory contexts, where a single concept may be referred to using multiple terminologies. By integrating synonym expansion into the BM25 framework, the system ensures comprehensive coverage and improves recall without sacrificing precision. Collectively, these enhancements enable the retrieval engine to support robust, flexible, and clinically meaningful harmonization of laboratory units.
@@ 2.2.5 Hybrid Lexical–Semantic Retrieval and Ranking
Our hybrid search system integrates both lexical and semantic search capabilities through a multi-tiered (Fig. 2) designed to enhance retrieval accuracy and contextual relevance. The first component of this architecture involves the generation of combined text embeddings that represent laboratory test descriptions alongside relevant metadata. These embeddings enable the system to capture nuanced semantic relationships between queries and candidate entries, going beyond surface-level term matching.
To refine retrieval precision, the system employs field-specific search boosts that reflect the relative importance of different attributes. For instance, fields such as standardized test names or unit designations may carry more weight in the matching process than auxiliary metadata. These boost factors are configurable and allow the system to dynamically adjust the influence of each field depending on the context of the search.
Structured queries are constructed to leverage both exact and fuzzy matching mechanisms, alongside semantic vector similarity. The system utilizes a custom-built module that parses incoming queries using field-aware syntax, applies the predefined boost factors, and computes a composite score for each candidate result. This score is derived through a weighted combination of BM25 lexical scores and cosine similarity from semantic embeddings, allowing for a balanced integration of symbolic and contextual retrieval. By merging these complementary strategies, the system achieves high recall and precision in laboratory unit harmonization tasks, even in the presence of terminological variation and incomplete data.
@@ 2.2.6 Tuning Lexical and Semantic Weights via Bayesian Optimization
To determine the optimal combination of retrieval methods, we implemented Bayesian optimization [13] to fine-tune the weights assigned to various components of the search process. This approach allows for the systematic exploration of the search parameter space, ensuring that the retrieval system operates at peak efficiency while balancing competing factors. The optimization process primarily focused on adjusting field boost weights, which control the relative importance of different fields (e.g., test descriptions or units) in the retrieval process. By tuning these parameters, the system can adapt to the specific needs of each search context.
Another key aspect of the optimization was balancing the contributions from the lexical (BM25) and semantic (embedding-based) search components. Both approaches offer unique advantages: lexical search excels at exact term matching, while semantic search captures deeper contextual relationships. The optimization process sought to find the best weighting between these two components to maximize retrieval effectiveness across diverse queries and data types.
Finally, the optimization aimed to maximize the Mean Reciprocal Rank (MRR) [14], a widely used performance metric that evaluates the ranking quality of retrieved results. By focusing on MRR, the system ensures that relevant harmonization candidates appear near the top of the result list, enhancing the user experience and enabling faster, more accurate decision-making in laboratory unit harmonization tasks. Through this optimization process, we were able to fine-tune the retrieval system to deliver the most relevant results while minimizing errors and inconsistencies.
@@ 2.3 Transformer-Based Reranking of Retrieved Candidates
The reranker module is responsible for ranking candidate harmonizations presented by previous modules by evaluating their semantic compatibility with the input laboratory test. It operates on test triads, each comprising a test name, a sample type, and a measurement unit. For each candidate triad, the reranker computes a compatibility score based on its similarity to the input, enabling the system to prioritize the most plausible harmonization suggestions. This component plays a critical role in filtering out false matches and refining the output quality of the harmonization pipeline.
To address the complexity of laboratory data, which often includes specialized terminology, institution-specific conventions, and heterogeneous representations, the reranker incorporates tokenization and handling mechanisms specifically adapted to biomedical language. This enables the model to more effectively understand and compare test descriptions that differ in surface form but refer to the same concept.
The reranker is further enhanced by contrastive learning objectives that help the model distinguish between closely related yet non-equivalent test triads. This is especially important for laboratory terms that may differ only slightly in wording or structure but convey distinct clinical meanings. By leveraging a pairwise comparison strategy, the reranker can evaluate subtle differences and improve ranking precision.
Importantly, the model is designed to support incremental learning. As the system is used in real-world settings, manually verified harmonizations are incorporated back into the training set. This continuous learning process allows the reranker to evolve over time and adapt to emerging terminology and use cases, improving its long-term performance and generalization capabilities.
@@ 2.4 System Performance Evaluation
We evaluated the performance of our approach through a comprehensive suite of metrics, ensuring that both the retrieval and re-ranking components of the system were thoroughly assessed. For retrieval performance, we measured metrics such as Precision@k, Recall@k, and Mean Reciprocal Rank (MRR). These metrics allow us to evaluate how well the system retrieves relevant candidates across different ranks, ensuring that the most relevant harmonization suggestions appear early in the result set and assessing the system’s overall retrieval effectiveness.
@@ 3.2 Sentence Embedding Selection
To select an appropriate sentence encoder for our retrieval task involving laboratory tests, samples, and unit combinations, we evaluated several pre-trained models based on their performance in the Massive Text Embedding Benchmark (MTEB), as shown in Table 1. Our primary selection criteria included model size—favoring models with fewer than one billion parameters for efficiency—and the overall mean task score, which reflects performance across a diverse set of semantic tasks. This balance ensured both strong generalization capabilities and practical resource requirements for integration into our system.
@@
Among the evaluated models meeting the size restriction, stella_en_400M_v5 — a transformer with approximately 435 million parameters trained using Matryoshka Representation Learning (MRL) [15] — demonstrated strong general performance. While this model produces 4096-dimensional embeddings, the Matryoshka training method distributes semantic information hierarchically across the embedding dimensions, allowing effective selection of lower-dimensional subsets without significant loss of information. Leveraging this property, we used 1024-dimensional representations for our experiments, balancing embedding compactness with semantic richness.
In comparison, lighter models such as bge-small-en-v1.5 and all-MiniLM-L6-v2, with 33 million and 22 million parameters respectively, showed lower Mean (Task) scores, reflecting a trade-off between model size and embedding quality.
Our approach uses these embedding models to generate a candidate set of matches based on semantic similarity. The responsibility of fine-grained selection according to the particularities of our application is delegated to a reranker module, which applies domain-specific criteria to reorder the candidates for the final retrieval.
We also explored fine-tuning using a synthetic dataset of sentence pairs labeled as semantically equivalent or dissimilar, focusing on laboratory-specific terminology. However, our experiments showed that general-purpose pre-trained models consistently outperformed these fine-tuned variants, indicating that large-scale general-domain training still provides strong out-of-the-box performance. Future work may revisit fine-tuning using alternative strategies or additional domain supervision.

### v3

@@ 1 Background and Significance
Modern clinical research increasingly relies on integrating heterogeneous data sources, yet inconsistent units of measurement remain a persistent challenge, particularly in laboratory analyses where identical quantities may be reported using diverse conventions. Harmonizing these units is essential for large-scale studies, enabling interoperability across electronic health record systems and supporting reliable, reproducible analyses. Traditional approaches often rely on manually curated rules and mapping tables developed by domain experts [1]. While effective in limited contexts, these methods are labor-intensive, difficult to scale, and require ongoing maintenance to accommodate new laboratory codes and units, introducing bottlenecks and delaying research workflows.
Recent advances in machine learning and information retrieval provide promising avenues for automated unit harmonization. Techniques such as sentence embeddings, bidirectional transformers, and optimized retrieval models have demonstrated improved scalability, adaptability, and contextual understanding [2, 3, 4, 5]. In this study, we present a novel automated unit harmonization system that integrates Bayesian-optimized BM25, sentence embeddings, and transformer-based re-ranking. This framework addresses major limitations of existing approaches, including poor scalability, limited adaptability to diverse naming conventions, and overreliance on manual validation, while incorporating dynamic feedback mechanisms.
Beyond enabling seamless data integration, harmonizing units has a direct impact on downstream clinical and research analytics. By consolidating columns representing the same measurement, redundancy is reduced, the number of observations per feature increases, and correlations between duplicate columns are eliminated. These improvements enhance statistical power, increase the reliability of feature importance scores, facilitate detection of associations, and reduce noise from inconsistent measurements. Consequently, unit harmonization strengthens the interpretability and robustness of analytical models, supporting more reliable, timely, and actionable scientific insights that can ultimately inform clinical decision-making.
Current approaches to unit harmonization span a spectrum from basic open-source tools to sophisticated commercial platforms. Open-source frameworks such as ehrapy [6], psHarmonize [7], and MIMIC-Extract [8] provide foundational harmonization capabilities but typically require substantial manual intervention and offer limited scalability for large datasets. Commercial platforms deliver advanced automation and can handle large-scale data processing, but their proprietary nature limits accessibility and customization for research applications.
Our proposed framework addresses this gap by combining Bayesian-optimized BM25, sentence embeddings, and transformer-based re-ranking to achieve advanced automation and scalability comparable to commercial systems while maintaining the flexibility and accessibility of open-source solutions. This hybrid approach enables the processing of billions of records with minimal manual intervention, providing both robustness and practical applicability for large-scale clinical research. Importantly, the system is designed with a modular architecture that makes it easily extendable to any alternative dataset, supporting diverse clinical coding standards including Loinc [9], ATC [10], MedDRA [11], RxNorm [12] and SNOMED-CT [13] beyond the unit harmonization focus of this study.
@@ 2.1.1 Primary Dataset
In this study, we used the 2024 Q1, Q2, and Q3 version of the Optum Clinformatics Data Mart. To support the harmonization process, key fields were extracted from the dataset, including LOINC codes, units of measurement, frequency of occurrence, and descriptive statistics such as minimum, maximum, mean, and standard deviation. This process resulted in approximately 30,000 unique triads (test, sample, unit) requiring standardization. Of these, 17,244 entries were identified in our internal database as candidates for matching.
@@ 2.2.1 Overall Pipeline Structure
The system is organized as an end-to-end processing pipeline (Fig. 1) comprising three main stages: preprocessing, harmonization, and post-processing. The workflow begins with preprocessing, which retrieves raw data from the source system, Optum Clinformatics Data Mart. This stage focuses on relevant elements such as LOINC codes—standardized identifiers for laboratory tests—and reported units. The data then undergoes cleaning and normalization to ensure consistency. A validation step follows, confirming the integrity of LOINC codes and associated metadata. Finally, the dataset is enriched by mapping LOINC codes to standardized labels and attributes, adding valuable information to support the subsequent harmonization phase.
@@
During harmonization, the system constructs tailored queries that generate ranked proposals for matching equivalents. These queries leverage the hybrid retrieval engine presented in this paper, enabling efficient searches for equivalent data elements. The hybrid retrieval engine returns candidate matches—potential harmonization targets retrieved from the reference database. These represent the initial pool of harmonization proposals returned by the system based on lexical (BM25) and semantic similarity scores, before refinement through the transformer-based reranking process. The retrieved results are then processed and structured for evaluation, followed by a reranking step that prioritizes candidates based on relevance scores to enhance overall accuracy.
@@ 2.2.2 Predictive Model Architecture
The Retrieval Reranker serves as the final decision layer. This bidirectional transformer model reevaluates and reorders candidate matches based on deeper contextual understanding, addressing subtle semantic distinctions critical in complex medical terminology.
@@ 2.2.3 Baseline Retrieval Using BM25
To enhance retrieval effectiveness, we incorporated curated synonym lists that expand the query space by mapping equivalent or related terms. These lists were manually developed by the primary author to address the extensive variability in unit and sample nomenclature found in clinical datasets. For laboratory units, the synonym mappings capture notation variants such as scientific notation formats (10*3/L, 10**3/L, 10^3/L), word-based expressions (THOUSAND/L, THOU/L), and symbolic alternatives (x10(3)/L, k/uL). For samples, they consolidate semantically equivalent terms such as “plasma, blood plasma, plas” and “serum, blood serum, ser,” while also resolving more complex combinations like “serum/plasma, serum or plasma, plasma/serum.” In total, the resource comprises approximately 300 unit-based and 50 sample-based synonym groups. These synonym mappings were implemented using Elasticsearch’s standard synonym filter, which integrates seamlessly with the BM25 framework by automatically expanding queries during indexing and search. Collectively, this strategy enables the retrieval engine to provide robust, flexible, and clinically meaningful harmonization of laboratory units.
@@ 2.2.5 Hybrid Lexical–Semantic Retrieval and Ranking
Our hybrid search system integrates both lexical and semantic search capabilities through a multi-tiered architecture (Fig. 2) designed to enhance retrieval accuracy and contextual relevance. The first component of this architecture involves the generation of combined text embeddings that represent laboratory test descriptions alongside relevant metadata. These embeddings enable the system to capture nuanced semantic relationships between queries and candidate entries, going beyond surface-level term matching.
To refine retrieval precision, the system employs field-specific search boosts that reflect the relative importance of different attributes. For instance, fields such as standardized test names or unit designations may carry more weight in the matching process than auxiliary metadata. These boost factors are configurable and allow the system to dynamically adjust the influence of each field depending on the context of the search. The main sources of error in the harmonization process stem from inconsistencies in three key components: test name, sample, and unit. These elements are essential for achieving correct matches, and discrepancies in any of them can lead to harmonization errors. To address this challenge, the model can be adjusted by optimizing the relative importance (weights) of each parameter for the specific training dataset, allowing better adaptation to dataset-specific characteristics and reducing potential mismatches.
Formally, the final ranking score for a candidate document $d$ given a query $q$ is computed as a weighted combination of lexical and semantic components:
$\text{Score}(d,q)=\alpha\cdot\text{BM25}(d,q)+\beta\cdot\max(0,\text{CosSim}(E_{q},E_{d}))$ (1)
where:
  • $\alpha$ and $\beta$ are optimized weights for lexical and semantic components, respectively,
  • $\text{BM25}(d,q)$ is the BM25 score as defined in the previous section,
  • $\text{CosSim}(E_{q},E_{d})$ is the cosine similarity between query embedding $E_{q}$ and document embedding $E_{d}$,
  • $\max(0,\cdot)$ clips negative similarities to zero, ensuring only positive semantic matches contribute.
For field-specific queries, the lexical component is further decomposed to explicitly account for the importance of harmonization-critical attributes:
$\text{BM25}(d,q)=\sum_{f\in\{\text{test},\text{sample},\text{unit},...\}}w_{f}\cdot\text{BM25}_{f}(d_{f},q_{f})$ (2)
where $w_{f}$ represents the weight for field $f$, and $\text{BM25}_{f}(d_{f},q_{f})$ is the BM25 score computed on the field-specific content. This formulation allows the system to explicitly emphasize the test, sample, and unit fields, which are the primary sources of harmonization errors.
Structured queries are constructed to leverage both exact and fuzzy matching mechanisms, alongside semantic vector similarity. A custom-built module parses incoming queries using field-aware syntax, applies the predefined boost factors, and computes a composite score according to the equations above. By merging symbolic (BM25) and contextual (semantic embeddings) strategies in a principled way, the system achieves high recall and precision in laboratory unit harmonization tasks, even in the presence of terminological variation and incomplete data.
@@ 2.2.6 Tuning Lexical and Semantic Weights via Bayesian Optimization
To determine the optimal combination of retrieval methods, we implemented Bayesian optimization to fine-tune the weights assigned to various components of the search process. This approach systematically explores the parameter space, ensuring that the retrieval system operates at peak efficiency while balancing competing factors. Specifically, the optimization process focuses on adjusting field boost weights, which control the relative importance of different fields (e.g., test descriptions, specimen types, and units) in the retrieval process. By tuning these parameters, the system can adapt to the specific requirements of each search context.
Formally, the Bayesian optimization seeks to identify the optimal parameter vector
$\theta=[\alpha,\beta,w_{\text{test}},w_{\text{sample}},w_{\text{unit}},$...]
that maximizes the Mean Reciprocal Rank (MRR) on a validation set:
$\theta^{*}=\arg\max_{\theta}\text{MRR}(\theta)=\arg\max_{\theta}\frac{1}{|Q|}\sum_{q\in Q}\frac{1}{\text{rank}_{q}(\theta)},$ (3)
where $Q$ is the set of validation queries and $\text{rank}_{q}(\theta)$ is the rank of the correct answer for query $q$ under parameter configuration $\theta$.
The optimization primarily targets the main sources of harmonization errors, which arise from inconsistencies in test names, sample types, and units. By systematically adjusting the weights of these critical parameters, the system minimizes mismatches and improves overall harmonization accuracy. This data-driven approach enables the retrieval model to adapt to dataset-specific characteristics, automatically learning the optimal balance between test name precision, sample specificity, and unit consistency.
In addition, the optimization balances the contributions of lexical (BM25) and semantic (embedding - based) components. Lexical search captures exact term matches, whereas semantic search captures contextual relationships. The weights of these components, $\alpha$ and $\beta$, are included in the parameter vector, allowing the system to learn an optimal combination that maximizes retrieval performance across diverse datasets.
The Bayesian optimization employs a Gaussian Process (GP) as a surrogate model with the Expected Improvement (EI) acquisition function:
$\alpha_{\text{EI}}(\theta)=\mathbb{E}[\max(0,f(\theta)-f(\theta^{+}))]$ (4)
where $f(\theta^{+})$ is the best observed MRR value so far, and the expectation is taken over the GP posterior. The optimization bounds are defined as:
$\alpha,\beta$ $\in[0,10]$ (5)
$w_{\text{test}},w_{\text{sample}},w_{\text{unit}}$ $\in[0,5]$ (6)
By focusing on MRR as the objective, the optimization ensures that relevant harmonization candidates appear near the top of the result list, thereby enhancing the user experience and enabling faster, more accurate decision-making in laboratory unit harmonization tasks.
@@ 2.3 Transformer-Based Reranking of Retrieved Candidates
The reranker module refines the ranking of candidate harmonizations returned by earlier retrieval stages by evaluating their semantic compatibility with the input laboratory test. Each laboratory entry is represented as a triad comprising a test name, a sample type, and a measurement unit. For each candidate triad, the reranker assigns a compatibility score that reflects its likelihood of being a valid harmonization, thereby prioritizing the most plausible candidates and filtering out false matches.
2.3.1 Transformer Architecture
The reranker is implemented using RoBERTa-large-PM-M3-Voc-hf (Figure 3), a domain-adapted variant of RoBERTa [19] pretrained on biomedical corpora including PubMed abstracts, PMC articles, and MIMIC-III clinical notes. The model comprises 24 transformer layers with a hidden dimension of 1024, 16 attention heads per layer, and a feed-forward dimension of 4096. Its specialized biomedical vocabulary contains 50,008 tokens, enabling effective tokenization of clinical terminology.
Figure 3: Detailed architecture of the RoBERTa-large-PM-M3-Voc-hf model showing the biomedical transformer components, training specifications, and domain adaptation features used in the reranker module.
For harmonization, we adapt the model by replacing the masked language modeling head with a binary classification head. The input format concatenates two triads—one from the query and one candidate—using the following template:
$\text{Input}=\texttt{<s>}\;T_{1}\;\texttt{</s></s>}\;T_{2}\;\texttt{</s>}$ (7)
where $T_{i}$ is a structured representation of the form:
$T_{i}=\texttt{TEST: \{test\_name\} SAMPLE: \{sample\_type\} UNIT: \{unit\}}$
Contextualized embeddings are produced through multi-head self-attention:
$\text{Attention}(Q,K,V)=\text{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$ (8)
with $Q$, $K$, and $V$ denoting the query, key, and value matrices for each head, and $d_{k}=64$ the dimension per head. The outputs of all heads are concatenated and linearly projected to produce the final token embeddings.
The embedding of the <s> token is used as the pooled sequence representation:
$h_{\texttt{<s>}}=\text{RoBERTa}(\text{Input})[:,0,:]$ (9)
Finally, the pooled representation is passed to a linear classification head, and the compatibility score is computed as:
$p_{\text{compatible}}=\sigma(W_{c}\cdot h_{\texttt{<s>}}+b_{c}),$ (10)
where $W_{c}$ and $b_{c}$ are learned parameters, and $\sigma$ is the sigmoid function for binary classification.
2.3.2 Contrastive Learning Objective
To train the reranker, we adopt a contrastive learning framework that teaches the model to differentiate semantically equivalent from non-equivalent triads. Positive examples are generated using curated biomedical synonym dictionaries, while negative examples are constructed through hierarchical corruption of one, two, or all three triad components. For the “Sample” and “Unit” components, we use the same synonym lists mentioned previously. For the test names, positive examples are created by randomly selecting among the different synonyms available in LOINC and our internal laboratory codes database.
Formally, the negative sets are defined as
$\mathcal{N}_{1}$ $=\{((t,s,u),(t^{\prime},s,u)):t^{\prime}\notin\text{syn}(t)\}$ (11)
$\mathcal{N}_{2}$ $=\{((t,s,u),(t^{\prime},s^{\prime},u)):t^{\prime}\notin\text{syn}(t),s^{\prime}\notin\text{syn}(s)\}$ (12)
$\mathcal{N}_{3}$ $=\{((t,s,u),(t^{\prime},s^{\prime},u^{\prime})):\text{all components differ}\},$ (13)
where $(t,s,u)$ denotes a source triad and $\text{syn}(\cdot)$ indicates synonym substitution according to the corresponding dictionaries or databases. This strategy ensures the model learns to recognize both subtle and substantial mismatches.
The training objective is binary cross-entropy with label smoothing:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right],$ (14)
where $y_{i}\in\{0,1\}$ is the compatibility label and $\hat{y}_{i}$ the predicted probability. The final training dataset comprised approximately 3.1 million balanced pairs across positive and negative categories.
To strengthen contrastive learning, we employed three strategies: (i) hard negative mining, selecting near-miss triads with overlapping terminology but different meanings; (ii) dynamic difficulty scheduling, progressively shifting the ratio of easy to hard negatives during training; and (iii) balanced sampling, ensuring equal representation of test name, sample, and unit mismatches.
2.3.3 Incremental Learning and Adaptation
The reranker supports incremental updates through continuous integration of manually verified harmonizations from real-world use. These feedback loops enable the model to adapt to emerging terminology, institution-specific conventions, and evolving laboratory practices. This incremental learning capability ensures long-term robustness and generalization in practical deployment scenarios.
@@ 2.4 System Performance Evaluation
The gold standard for evaluation consisted of 2,500 randomly selected triads (TEST, SAMPLE, UNIT) manually annotated by the primary author. We evaluated the performance of our approach through a comprehensive suite of metrics, ensuring that both the retrieval and re-ranking components of the system were thoroughly assessed. For retrieval performance, we measured metrics such as Precision@k, Recall@k, and Mean Reciprocal Rank (MRR). These metrics allow us to evaluate how well the system retrieves relevant candidates across different ranks, ensuring that the most relevant harmonization suggestions appear early in the result set and assessing the system’s overall retrieval effectiveness.
@@ 3.2 Sentence Embedding Selection
We selected candidate pre-trained sentence encoders for our retrieval task involving laboratory tests, samples, and unit combinations, guided by their reported performance on the Massive Text Embedding Benchmark (MTEB) [20] (Table 1). Selection prioritized models with fewer than one billion parameters for efficiency and high Mean (Task) scores, reflecting robust general-purpose semantic performance. By emphasizing generalization over task-specific optimization, the chosen model can be reused across diverse retrieval or semantic tasks, ensuring scalability and versatility beyond this particular application.
@@
Among models satisfying our size constraint, stella_en_400M_v5—a 435M-parameter transformer trained with Matryoshka Representation Learning (MRL) [21]—offered the best balance between performance and efficiency. Although it produces 4096-dimensional embeddings, MRL distributes semantic information hierarchically, allowing effective use of 1024-dimensional subsets without significant loss of information. Smaller models, such as bge-small-en-v1.5 and all-MiniLM-L6-v2, achieved lower Mean (Task) scores, illustrating the trade-off between model size and embedding quality.
These embeddings are used to generate candidate matches based on semantic similarity, with a reranker module applying domain-specific criteria for final selection. In our initial experiments, fine-tuning on a synthetic laboratory-specific dataset did not outperform the general-purpose pre-trained embeddings, highlighting the advantage of large-scale general-domain training for robust and reusable sentence representations.
It is important to note that this model selection was performed at the time of the study, and future research may identify improved models of similar size that could offer better performance, making this an aspect that can and should be revisited as the field evolves.